# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v2

@@ 6 Structured bandits with an Infinite Function Class
This result mirrors Theorem 3.3, with the function gap replaced by $\varepsilon$, allowing for instance-dependent optimization of $\varepsilon$ and tighter bounds. The proof for part (a) carries over with simple modifications. In contrast, proving part (b) is considerably more subtle. In the infinite case, $\mathtt{Greedy}$ may not get stuck on a single reward function—it could almost surely switch among infinitely many. The key insight is that such fluctuations need not impact the arm choice: even as the predictor changes, the greedy selection may remain fixed. The proof exploits this decoupling, constructing events where the algorithm persistently selects a decoy arm, even as the greedy predictors continue to evolve.
@@ 7 Examples
Let us instantiate our characterization for several well-studied reward structures from bandits literature. We consider linear, Lipschitz, and (one-dimensional) polynomial structures, for bandits as well as contextual bandits. All reward structures in this section are discretized to ensure finiteness, as required for our complete characterization in Sections 3, 4 and 5. (While our partial characterization in Section 6 handles infinite reward structures, a secondary goal of this section is to illustrate how common infinite reward structures can be meaningfully discretized so that the complete finite-structure results become directly applicable.) The discretization is consistent across different reward functions, in the sense that all functions take values in the same (discrete) set $\mathcal{R}$, with $|\mathcal{R}|\ll|\mathcal{F}|$. This prevents a trivial form of self-identifiability that could arise if each reward function $f$ were discretized independently and inconsistently, resulting in some $f(a)$ values being unique and making $f$ self-identifiable solely due to the discretization strategy specific to $f$ [^15] For example, consider an instance $(f^{*},\mathcal{F})$ being not self-identifiable, with its decoy $f_{\mathtt{dec}}\in\mathcal{F}$ satisfying $f^{*}(a_{\mathtt{dec}})=f_{\mathtt{dec}}(a_{\mathtt{dec}})=0.5$. Now, suppose we discretize $f^{*}(a_{\mathtt{dec}})$ using discretization step 0.1 and discretize $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ using discretization step 0.2. After this modification, $f^{*}(a_{\mathtt{dec}})$ and $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ would no longer be equal, and self-identifiability could occur.
@@
In what follows, we present each structure in a self-contained way, interpreting it as special case of our framework. Since our presentation focuses on best-arm-unique reward functions, our examples are focused similarly (except those for Linear CB). Throughout, let $[y,y^{\prime}]_{\varepsilon}$ be a uniform discretization of the $[y,y^{\prime}]$ interval with step $\varepsilon>0$, namely: $[y,y^{\prime}]_{\varepsilon}:=\left\{\,\varepsilon\cdot n\in[y,y^{\prime}]:\,n\in\mathbb{N}\,\right\}$. Likewise, we define $(y,y^{\prime})_{\varepsilon}:=\left\{\,\varepsilon\cdot n\in(y,y^{\prime}):\,n\in\mathbb{N}\,\right\}$.
@@
Linear bandits is a well-studied variant of bandits (Auer, 2002; Abe et al., 2003; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). [^16] We consider stochastic linear bandits. A more general model of adversarial linear bandits is studied since Awerbuch and Kleinberg (2008); McMahan and Blum (2004), see Bubeck and Cesa-Bianchi (2012, Chapter 5) for a survey. Formally, it is a special case of $\mathtt{StructuredMAB}$ defined as follows. Arms are real-valued vectors: $\mathcal{A}\subset\mathbb{R}^{d}$, where $d\in\mathbb{N}$ is the dimension. Reward functions are given by $f_{\theta}(a)=a\cdot\theta$ for all arms $a$, where $\theta\in\Theta\subset\mathbb{R}^{d}$. The parameter set $\Theta$ is known to the algorithm, so the function class is =$\mathcal{F}=\left\{\,f_{\theta}:\,\theta\in\Theta\,\right\}$. The true reward function is $f^{*}=f_{\theta^{*}}$ for some $\theta^{*}\in\Theta$. (Fixing $\Theta$, we interpret $\theta^{*}$ as a “problem instance”.)
Linear bandits, as traditionally defined, let $\Theta$ be (continuously) infinite, e.g., a unit $\ell_{1}$-ball, and sometimes consider an infinite (namely, convex) action set. Here, we consider a “discretized” version, whereby both $\Theta$ and $\mathcal{A}$ are finite. Specifically, $\Theta=\left(\,[-1,1]_{\varepsilon}\setminus\{0\}\,\right)^{d}$, i.e., all parameter vectors in $[0,1]^{d}$ with discretized non-zero coordinates. Action set $\mathcal{A}$ is an arbitrary finite subset of $[-1,1]^{d}$ containing the hypercube $\left\{\,-1,1\,\right\}^{d}$.  [^17] For ease of exposition, we relax the requirement that expected rewards must lie in $[0,1]$. Note that each reward function $f_{\theta}$, $\theta\in\Theta$ has a unique best arm $a^{*}_{\theta}=\mathrm{sign}(\theta):=\left(\,\mathrm{sign}(\theta_{i}):\,i\in[d]\,\right)\in\left\{\,-1,1\,\right\}^{d}$.
@@
Let $j\in[d]$ be a coordinate with the smallest $|\theta^{*}_{j}|$. Choose arm $a_{\mathtt{dec}}\in\left\{\,-1,1\,\right\}^{d}$ with $(a_{\mathtt{dec}})_{i}=\mathrm{sign}(\theta^{*}_{i})$ for all coordinates $i\neq j$, and flipping the sign for $i=j$, $(a_{\mathtt{dec}})_{j}=-\mathrm{sign}(\theta^{*}_{j})$. Note that $\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}=\|\theta^{*}\|_{1}-2\,\min_{i\in[d]}|\theta^{*}_{i}|\in[d\varepsilon,d]$.
Now, for any given $\alpha\in[d\varepsilon,\,d]_{\varepsilon}$ and any sign vector $v\in\{-1,1\}^{d}$, there is $\theta\in\Theta$ such that $\|\theta\|_{1}=\alpha$ and its signs are aligned as $\mathrm{sign}(\theta)=v$. Thus, there exists $\theta_{\mathtt{dec}}\in\Theta$ such that $\|\theta_{\mathtt{dec}}\|_{1}=\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}$ and $\mathrm{sign}(\theta_{\mathtt{dec}})=a_{\mathtt{dec}}$. Note that $a_{\mathtt{dec}}$ is the best arm for $\theta_{\mathtt{dec}}$. Moreover, $\innerproduct{\theta_{\mathtt{dec}}}{a_{\mathtt{dec}}}=\|\theta_{\mathtt{dec}}\|_{1}=\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}<\|\theta^{*}\|_{1}=\innerproduct{\theta^{*}}{a^{*}}$. So, $\theta_{\mathtt{dec}}$ is a decoy for $\theta^{*}$.
@@
Like in linear bandits, the function class is =$\mathcal{F}=\left\{\,f_{\theta}:\,\theta\in\Theta\,\right\}$. The true reward function is $f^{*}=f_{\theta^{*}}$ for some $\theta^{*}\in\Theta$, which we interpret as a “problem instance”.
@@
[…]
@@
Consider vector $\theta_{\mathtt{dec}}\in\Theta$ such that it coincides with $\theta^{*}$ on the first $d-1$ components, and $(\theta_{\mathtt{dec}})_{d}=-1$. We claim that $\theta_{\mathtt{dec}}$ is a decoy for $\theta^{*}$.
To prove this claim, fix context $x\in\mathcal{X}$. Let $a^{*},a_{\mathtt{dec}}$ be some optimal arms for this context under $\theta^{*}$ and $\theta_{\mathtt{dec}}$, respectively. Then $a_{\mathtt{dec}}\in[K-1]$. (This is because the expected reward $x(a)\cdot\theta^{*}$ of arm $a$ is greater than -1 when $a\in[K-1]$, and exactly $-1$ when $a=K$.) Similarly, we show that $a^{*}=K$. It follows that $x(a_{\mathtt{dec}})\cdot\theta_{\mathtt{dec}}=x(a_{\mathtt{dec}})\cdot\theta^{*}$, since $\theta_{\mathtt{dec}}$ and $\theta^{*}$ coincide on the first $K-1$ coordinates, and the last coordinate of $x(a_{\mathtt{dec}})$ is 0. Moreover $x(a_{\mathtt{dec}})\cdot\theta^{*}<1=x(a^{*})\cdot\theta^{*}$. Putting this together, $x(a_{\mathtt{dec}})\cdot\theta_{\mathtt{dec}}=x(a_{\mathtt{dec}})\cdot\theta^{*}<x(a^{*})\cdot\theta^{*}$, completing the proof.
@@
[…]
@@
[…]
@@
We define polynomial bandits as a special case of $\mathtt{StructuredMAB}$ with action set $\mathcal{A}\subset\mathbb{R}$ and reward functions $f$ are degree-$p$ polynomials, for some degree $p\in\mathbb{N}$. Denote reward functions as $f=f_{\bm{\theta}}$, where ${\bm{\theta}}=(\theta_{0},\dots,\theta_{p})\in\mathbb{R}^{p+1}$ is the parameter vector with $\theta_{p}\neq 0$, so that $f_{\bm{\theta}}(a)=\sum_{q=0}^{p}\theta_{q}\cdot a^{q}$ for all arms $a$. The function set is =$\mathcal{F}=\left\{\,f_{\bm{\theta}}:\,{\bm{\theta}}\in\Theta\,\right\}$, for some parameter set $\Theta$. Typically one allows continuously many actions and parameters, i.e., an infinite reward structure.
@@
[…]
@@
Consider discretized quadratic bandits, for some fixed discretization step $\varepsilon\in(0,\nicefrac{{1}}{{2}}]$. Fix any leading coefficient $\gamma\in[-1,-0.5]_{\varepsilon}$. Then for any reward function $f^{*}=f_{(\gamma,\mu,c)}\in\mathcal{F}_{\gamma}$, it has a decoy $f_{\mathtt{dec}}\in\mathcal{F}_{\gamma}$, as long as $\mu,c$ are bounded away from 00: $\mu\geq\varepsilon$ and $c\geq|\gamma|\varepsilon^{2}$.
@@ 8 Self-identifiability makes the problem easy
$\mathcal{E}_{\mathtt{MAB}}$ :={|$:=\left\{\,\absolutevalue{\bar{r}_{t}(a)-f^{*}(a)}>\beta_{t}\left(\,N_{t}(a)\,\right)\quad\forall a\in\mathcal{A},\,t\in\mathbb{N}\,\right\},$ (8.1)
$\mathcal{E}_{\mathtt{CB}}$ :={|$:=\{\;\absolutevalue{\bar{r}_{t}(x,a)-f^{*}(x,a)}<\beta_{t}\left(\,N_{t}(x,a)\,\right)\text{ and }N_{t}(x,a)>\Omega(N_{t}(\pi)\;p_{0})$
@@
|$\absolutevalue{\bar{r}_{t}(x,\pi(x))-f^{*}(x,\pi(x))}<\beta(T^{\prime})<\Gamma(f^{*}).$
Then recall for any function $f$ and context-arm pair $(x,a)$, we have either $f(x,a)=f^{*}(x,a)$ or |$\absolutevalue{f(x,a)-f^{*}(x,a)}\geq\Gamma(f^{*})$. This precisely means the policy $\pi$ becomes identified, and by self-identifiability, any information-aware algorithm will not keep choosing $\pi$. Hence, the total regret of the information-aware algorithm is at most $O(T^{\prime}|\Pi|)$, which is sublinear $o(T)$.
@@ 9 Conclusions
We study the greedy algorithm in structured bandits and characterize its asymptotic success vs failure in terms of a simple partial-identifiability property of the problem structure. Our characterization holds for arbitrary finite structures and extends to bandits with contexts and/or auxiliary feedback. In particular, we find that the greedy algorithm succeeds only if the problem is intrinsically “easy” for any algorithm which satisfies a mild non-degeneracy condition. We also provide a partial characterization for $\mathtt{StructuredMAB}$ with infinite reward structures (and finite action sets).
We provide several examples, both positive and negative, where we instantiate our characterization for various reward structures studied in the literature. We find that failure tends to be a common case for bandits (under commonly studied reward structures), whereas both failure and success are common for structured contextual bandits.
We identify three directions for further work. First, extend our positive and negative results to infinite action sets (and infinite function/model classes). Ideally one would like to obtain a full characterization, like in the finite case. Second, extend our characterization to approximate greedy algorithms, stemming either from approximate regression or from human behaviorial biases. A class of approximately-greedy algorithms, representing myopic human behavior under behavioral biases, was studied in Banihashem et al. (2023a, b), but only for unstructured multi-armed bandits. Third, while our “asymptotic” perspective enables a sharp characterization over reward structures, stronger guarantees are desirable for a particular reward structure. More concretely: negative results with failure probabilities guaranteed to be not extremely small, and/or positive results with regret rates that are guaranteed to not have extremely large “constant factors”. Such guarantees are only known for a few specific reward structures. [^21] Specifically, prior work provides strong positive guarantees for linear contextual bandits with smoothed/diverse contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023), and strong negative guarantees for unstructured multi-armed bandits and Bayesian bandits with priors independent across arms (Banihashem et al., 2023a, b).

### v3

@@ 6 Structured bandits with an Infinite Function Class
This result mirrors Theorem 3.3, with the function gap replaced by $\varepsilon$, allowing for instance-dependent optimization of $\varepsilon$ and tighter bounds. The proof for part (a) carries over with simple modifications. In contrast, proving part (b) is considerably more subtle. In the infinite case, $\mathtt{Greedy}$ may not get stuck on a single reward function—it could almost surely switch among infinitely many. The key insight is that such fluctuations need not impact the arm choice: even as the predictor $f_{t}$ changes, the greedy selection $a_{t}$ may remain fixed. The proof exploits this decoupling, constructing events where the algorithm persistently selects a decoy arm, even as the greedy predictors continue to evolve.
Discussion: challenges. An “if and only if” characterization for arbitrary infinite function classes is very difficult. First, one can no longer rely on the function-gap being strictly positive, which is a cornerstone of our analysis in the finite case. Second, $\mathtt{Greedy}$’s behavior can be highly unstable: the algorithm’s predictor $f_{t}$ can fluctuate indefinitely within a continuous region of functions that are all similarly consistent with data yet induce very different greedy action choices. As a result, the intuitive logic of “getting stuck in a decoy and staying there forever” does not directly extend.
The partial characterization in Theorem 6.2 is our proposed route to address these challenges. The margin $\varepsilon$ serves a dual purpose: it stands in for the now-absent function-gap, and it allows us to deal with the predictor’s instability (by showing that $a_{t}$ can remain permanently fixed).
However, this is still insufficient for a complete characterization. For many natural function classes, our framework leaves a set of instances, typically of fraction $O(\varepsilon)$, uncharacterized. The boundary between success and failure instances in a general infinite space can be highly complex; success instances can be very close to failure instances, making a sharp separation difficult. Our $\varepsilon$-interior notion is designed to provide a robust buffer around this boundary, at the cost of leaving instances within that buffer “undecided.” A tight characterization in full generality would likely require a more fine-grained analysis exploiting additional structural properties of the function class $\mathcal{F}$.
@@ 7 Examples
Let us instantiate our characterization for several well-studied reward structures from bandits literature. We consider linear, Lipschitz, and (one-dimensional) polynomial structures, for bandits as well as contextual bandits. All reward structures in this section are discretized to ensure finiteness, as required for our complete characterization in Sections 3, 4 and 5. (While our partial characterization in Section 6 handles infinite reward structures, a secondary goal of this section is to illustrate how common infinite reward structures can be meaningfully discretized so that the complete finite-structure results become directly applicable.) The discretization is consistent across different reward functions, in the sense that all functions take values in the same (discrete) set $\mathcal{R}$, with $|\mathcal{R}|\ll|\mathcal{F}|$. This prevents a trivial form of self-identifiability that could arise if each reward function $f$ were discretized independently and inconsistently, resulting in some $f(a)$ values being unique and making $f$ self-identifiable solely due to the discretization strategy specific to ff. [^15] For example, consider an instance $(f^{*},\mathcal{F})$ being not self-identifiable, with its decoy $f_{\mathtt{dec}}\in\mathcal{F}$ satisfying $f^{*}(a_{\mathtt{dec}})=f_{\mathtt{dec}}(a_{\mathtt{dec}})=0.5$. Now, suppose we discretize $f^{*}(a_{\mathtt{dec}})$ using discretization step 0.1 and discretize $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ using discretization step 0.2. After this modification, $f^{*}(a_{\mathtt{dec}})$ and $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ would no longer be equal, and self-identifiability could occur.
@@
In what follows, we present each structure in a self-contained way, interpreting it as special case of our framework. Since our presentation focuses on best-arm-unique reward functions, our examples are focused similarly (except those for Linear CB). Throughout, let $y,y^{\prime}]_{\varepsilon}$ be a uniform discretization of the $y,y^{\prime}]$ interval with step $\varepsilon>0$, namely: $y,y^{\prime}]_{\varepsilon}:=\left\{\,\varepsilon\cdot n\in[y,y^{\prime}]:\,n\in\mathbb{N}\,\right\}$. Likewise, we define $(y,y^{\prime})_{\varepsilon}:=\left\{\,\varepsilon\cdot n\in(y,y^{\prime}):\,n\in\mathbb{N}\,\right\}$.
@@
Linear bandits is a well-studied variant of bandits (Auer, 2002; Abe et al., 2003; Dani et al., 2008; Rusmevichientong and Tsitsiklis, 2010). [^16] We consider stochastic linear bandits. A more general model of adversarial linear bandits is studied since Awerbuch and Kleinberg (2008); McMahan and Blum (2004), see Bubeck and Cesa-Bianchi (2012, Chapter 5) for a survey. Formally, it is a special case of $\mathtt{StructuredMAB}$ defined as follows. Arms are real-valued vectors: $\mathcal{A}\subset\mathbb{R}^{d}$, where $d\in\mathbb{N}$ is the dimension. Reward functions are given by $f_{\theta}(a)=a\cdot\theta$ for all arms $a$, where $\theta\in\Theta\subset\mathbb{R}^{d}$. The parameter set $\Theta$ is known to the algorithm, so the function class is $\mathcal{F}=\left\{\,f_{\theta}:\,\theta\in\Theta\,\right\}$. The true reward function is $f^{*}=f_{\theta^{*}}$ for some $\theta^{*}\in\Theta$. (Fixing $\Theta$, we interpret $\theta^{*}$ as a “problem instance”.)
Linear bandits, as traditionally defined, let $\Theta$ be (continuously) infinite, e.g., a unit $\ell_{1}$-ball, and sometimes consider an infinite (namely, convex) action set. Here, we consider a “discretized” version, whereby both $\Theta$ and $\mathcal{A}$ are finite. Specifically, $\Theta=\left(\,[-1,1]_{\varepsilon}\setminus\{0\}\,\right)^{d}$, i.e., all parameter vectors in $0,1]^{d}$ with discretized non-zero coordinates. Action set $\mathcal{A}$ is an arbitrary finite subset of $1,1]^{d}$ containing the hypercube $\left\{\,-1,1\,\right\}^{d}$.  [^17] For ease of exposition, we relax the requirement that expected rewards must lie in $[0,1]$. Note that each reward function $f_{\theta}$, $\theta\in\Theta$ has a unique best arm $a^{*}_{\theta}=\mathrm{sign}(\theta):=\left(\,\mathrm{sign}(\theta_{i}):\,i\in[d]\,\right)\in\left\{\,-1,1\,\right\}^{d}$.
@@
[…]
@@
Like in linear bandits, the function class is $\mathcal{F}=\left\{\,f_{\theta}:\,\theta\in\Theta\,\right\}$. The true reward function is $f^{*}=f_{\theta^{*}}$ for some $\theta^{*}\in\Theta$, which we interpret as a “problem instance”.
@@
Consider linear CB with parameter set $\Theta=[-1,1]^{d}_{\varepsilon}$, for some degree $d\geq 2$ and discretization step $\varepsilon\in(0,\nicefrac{{1}}{{2}}]$ with $1/\varepsilon\in\mathbb{N}$. Suppose the context set is $\mathcal{X}=\prod_{a\in\mathcal{A}}S_{a}$, where $S_{a}\subset[-1,1]^{d}$ are the “per-arm” context sets. Assume $\mathrm{span}(S_{1},\dots,S_{K-1})\subset R^{d-1}$ and $S_{K}=\left\{\left(\,0,0,\dots,0,1\,\right)\right\}$. Then any instance $\theta^{*}\in\Theta$ with $\theta^{*}_{d}=1$ and $\norm{\theta^{*}}_{1}<2$ has a decoy in $\Theta$.
@@
Consider vector $\theta_{\mathtt{dec}}\in\Theta$ such that it coincides with $\theta^{*}$ on the first d1d-1 components, and $(\theta_{\mathtt{dec}})_{d}=-1$. We claim that $\theta_{\mathtt{dec}}$ is a decoy for $\theta^{*}$.
To prove this claim, fix context $x\in\mathcal{X}$. Let $a^{*},a_{\mathtt{dec}}$ be some optimal arms for this context under $\theta^{*}$ and $\theta_{\mathtt{dec}}$, respectively. Then $a_{\mathtt{dec}}\in[K-1]$. (This is because the expected reward $x(a)\cdot\theta^{*}$ of arm $a$ is greater than -1 when $a\in[K-1]$, and exactly 1-1 when $a=K$.) Similarly, we show that $a^{*}=K$. It follows that $x(a_{\mathtt{dec}})\cdot\theta_{\mathtt{dec}}=x(a_{\mathtt{dec}})\cdot\theta^{*}$, since $\theta_{\mathtt{dec}}$ and $\theta^{*}$ coincide on the first K1K-1 coordinates, and the last coordinate of $x(a_{\mathtt{dec}})$ is 0. Moreover $x(a_{\mathtt{dec}})\cdot\theta^{*}<1=x(a^{*})\cdot\theta^{*}$. Putting this together, $x(a_{\mathtt{dec}})\cdot\theta_{\mathtt{dec}}=x(a_{\mathtt{dec}})\cdot\theta^{*}<x(a^{*})\cdot\theta^{*}$, completing the proof.
@@
Define reward function $f_{\mathtt{dec}}$ by $f_{\mathtt{dec}}(a^{\prime})=\min\left(\,0,\;f(a)-\mathcal{D}(a,a^{\prime})\,\right)$ for all arms $a^{\prime}\in\mathcal{A}$. So, $f_{\mathtt{dec}}$ takes values in $\mathcal{R}$ and is Lipschitz w.r.t. $\mathcal{D}$ (since $\mathcal{D}$ satisfies triangle inequality); hence $f_{\mathtt{dec}}\in\mathcal{F}$. Also, $f_{\mathtt{dec}}$ has a unique best arm $a$ (since $f(a)>0$ and the distance between any two distinct points is positive). Note that $f_{\mathtt{dec}}(a)=f(a)<f(a^{*})$, so $f_{\mathtt{dec}}$ is a decoy.
@@
Define reward function $f_{\mathtt{dec}}$ by $f_{\mathtt{dec}}(x,a)=\min\left(\,0,\;f(x,\pi(x))-\mathcal{D}\left(\,(x,\pi(x)),\,(x,a)\,\right)\,\right)$ for all context-arm pairs $(x,a)$. Like in the proof of Lemma 7.9, we see that $f_{\mathtt{dec}}$ takes values in $\mathcal{R}$ and is Lipschitz w.r.t. $\mathcal{D}$, hence $f_{\mathtt{dec}}\in\mathcal{F}$. And it has a unique best arm $\pi(x)$ for each context $x$. Finally, $f_{\mathtt{dec}}(x,\pi(x))=f(x,\pi(x))<f(x,\,\pi^{*}(x))$, so $f_{\mathtt{dec}}$ is a decoy.
@@
We define polynomial bandits as a special case of $\mathtt{StructuredMAB}$ with action set $\mathcal{A}\subset\mathbb{R}$ and reward functions $f$ are degree-$p$ polynomials, for some degree $p\in\mathbb{N}$. Denote reward functions as $f=f_{\bm{\theta}}$, where ${\bm{\theta}}=(\theta_{0},\dots,\theta_{p})\in\mathbb{R}^{p+1}$ is the parameter vector with $\theta_{p}\neq 0$, so that $f_{\bm{\theta}}(a)=\sum_{q=0}^{p}\theta_{q}\cdot a^{q}$ for all arms $a$. The function set is $\mathcal{F}=\left\{\,f_{\bm{\theta}}:\,{\bm{\theta}}\in\Theta\,\right\}$, for some parameter set $\Theta$. Typically one allows continuously many actions and parameters, i.e., an infinite reward structure.
@@
Second, we claim that ${\bm{\theta}}_{\mathtt{dec}}\in\Theta_{\gamma}$. Indeed, the above equations imply that all coefficients of ${\bm{\theta}}_{\mathtt{dec}}$ are suitably discretized: $({\bm{\theta}}_{\mathtt{dec}})_{q}\in(-\infty,\infty)_{\delta(q)}$ for all $q\in\left\{\,0\,,\ \ldots\ ,p-1\,\right\}$. It remains to show that they are suitably bounded; this is where we use ${\bm{\theta}}\in\Theta^{\mathtt{bdd}}$. We argue this as follows:
@@
Consider discretized quadratic bandits, for some fixed discretization step $\varepsilon\in(0,\nicefrac{{1}}{{2}}]$. Fix any leading coefficient $\gamma\in[-1,-0.5]_{\varepsilon}$. Then for any reward function $f^{*}=f_{(\gamma,\mu,c)}\in\mathcal{F}_{\gamma}$, it has a decoy $f_{\mathtt{dec}}\in\mathcal{F}_{\gamma}$, as long as $\mu,c$ are bounded away from 0: $\mu\geq\varepsilon$ and $c\geq|\gamma|\varepsilon^{2}$.
@@ 8 Self-identifiability makes the problem easy
[…]
@@
[…]
@@ 9 Conclusions
We study $\mathtt{Greedy}$ in structured bandits and characterize its asymptotic success vs failure in terms of a simple partial-identifiability property of the problem structure. Our characterization holds for arbitrary finite structures and extends to bandits with contexts and/or auxiliary feedback. In particular, we find that $\mathtt{Greedy}$ succeeds only if the problem is intrinsically “easy” for any algorithm which satisfies a mild non-degeneracy condition. We also provide a partial characterization for $\mathtt{StructuredMAB}$ with infinite reward structures (and finite action sets).
Several examples, both positive and negative, instantiate our characterization for various well-studied reward structures. We find that failure tends to be a common case for bandits, whereas both failure and success are common for structured contextual bandits.
We identify three directions for further work. First, extend our characterization to infinite action sets and infinite function/model classes (ideally with a complete characterization, as discussed in Section 6). Second, consider approximate greedy algorithms, stemming either from approximate regression or from human behaviorial biases. Such algorithms, representing myopic human behavior under behavioral biases, were studied in Banihashem et al. (2023), but only for unstructured multi-armed bandits. Third, while our “asymptotic” perspective enables a general characterization, stronger regret guarantees are desirable for particular reward structures (and are known for only a few).